# PropMolFlow: Property-guided Molecule Generation with Geometry-Complete Flow Matching

### v3

@@ Abstract
Molecule generation is advancing rapidly in chemical discovery and drug design. Flow matching methods have recently set the state of the art (SOTA) in unconditional molecule generation, surpassing score-based diffusion models. However, diffusion models still lead in property-guided generation. In this work, we introduce PropMolFlow, a novel approach for property-guided molecule generation based on geometry-complete SE(3)-equivariant flow matching. Integrating five different property embedding methods with a Gaussian expansion of scalar properties, PropMolFlow achieves competitive performance against previous SOTA diffusion models in conditional molecule generation across various properties while preserving the stability and validity of the generated molecules, consistent with its unconditional counterpart. Additionally, it enables faster sampling speed with significantly fewer time steps compared to baseline models. We highlight the importance of validating the properties of generated molecules through DFT calculations performed at the same level of theory as the training data. Specifically, our analysis identifies properties that require DFT validation and others where a pretrained SE(3) geometric vector perceptron regressors provide sufficiently accurate predictions on generated molecules. Furthermore, we introduce a new task to assess the model’s ability to propose molecules with underrepresented property values, assessing its capacity for out-of-distribution generalization. Our findings reveal shortcomings in existing structural metrics, which mistakenly validate open-shell molecules or molecules with invalid valence-charge configurations, underscoring the need for improved evaluation frameworks. Overall, this work paves the way for developing targeted property-guided generation methods, enhancing the design of molecular generative models for diverse applications.
Keywords Molecule Generation  $\cdot$ Property Guidance  $\cdot$ Flow Matching  $\cdot$ SE(3) Equivariant Graph Neural Networks
@@ Introduction
Deep generative models show great promise to accelerate chemical discovery and drug design by direct statistical sampling of stable structures compared to resource-intensive physics-based approaches [1, 2]. The past few years have witnessed rapid development of generative methods for molecule generation, such as variational autoencoders (VAEs) [3, 4], generative adversarial networks (GANs) [5], autoregressive models [6], score-based diffusion models [7, 8], flow-based models [9, 10, 11] and language models [12, 13]. The predominant methods for molecule generation in 3D apply diffusion models combined with equivariant graph neural networks to generate molecules directly in 3D, providing a more accurate representation of molecular configurations compared to 1D SMILES and 2D graph-based representations [7, 14, 11, 8]. Flow matching, in particular, has emerged as a flexible alternative to diffusion models, providing greater control over the choice of prior distributions and probability paths, as well as enhanced sampling efficiency [15, 16, 17]. Its advantages have led to improved performance across various applications, including materials generation [18], protein design [19, 20], and molecule generation [21].
Seminal examples of diffusion models and flow matching for molecule generation treated molecules as point clouds and adopted a continuous representation of the atom types, leveraging E(3)-equivariant graph neural networks (EGNN) to learn the score function in diffusion models and the conditional velocity field in flow matching [7, 21]. However, the continuous representation of discrete molecular features, such as atom types and bond orders, fails to preserve their inherent discreteness, thus compromising their fidelity. Additionally, the EGNN architecture fails to capture molecular chirality, as an SE(3) generative process is required to account for chiral centers [22]. Recent studies also highlight the importance of incorporating bond orders in molecular graphs to improve the stability and validity of generated molecules [23, 24]. Building on these insights, recent diffusion models and flow matching approaches adopt a geometry-complete SE(3) generative process that incorporates bond orders into molecular graph representations, and utilize discrete diffusion or discrete flow matching for discrete molecular features such as atom types and bond orders [25, 26]. Notably, FlowMol integrates these advances to achieve SOTA performance in unconditional molecule generation [26]. However, its formulation remains limited to the unconditional setting, leaving its extension to property-guided molecule generation unaddressed.
@@
Conditional molecule generation involves designing new molecules that satisfy specific constraints or target properties. This is typically implemented by concatenating property values with node features to embed property information into molecular graphs, a technique that has been primarily demonstrated on the QM9 dataset [7, 8, 27, 25, 28]. While this approach has proved effective, it may oversimplify the interaction between molecular graphs and properties. For example, Gebauer et al. suggest that applying a Gaussian expansion to a scalar property may help extract property information more robustly [29]. Systematic studies of property embedding methods are lacking, and optimal embedding strategies may vary depending on the specific property. Moreover, previous methods use a pretrained classifier as the property predictor. Since this classifier is trained on the QM9 dataset, where all molecules are fully relaxed, it may struggle to make accurate predictions for generated molecules, which are likely not relaxed. Consequently, the reported results can be unreliable due to the limited generalization capacity of the pretrained classifiers/regressors used for evaluation. Last but not least, property-guided generation is typically evaluated in an ‘in-distribution’ setting, where property values sampled from within the QM9 distribution are used to generate molecules expected to exhibit the corresponding property values. In practice, for molecular discovery, we may want to evaluate the model’s ability to push the envelope of the original training distribution and generate molecules with underrepresented property values.
In this work, we introduce PropMolFlow—a novel property-guided molecule generation framework—that integrates various property embedding methods with an SOTA SE(3)-equivariant flow matching process based on geometric vector perceptrons (GVP). We evaluate PropMolFlow on the QM9 dataset, showing that it outperforms previous diffusion models and flow matching methods, setting a new SOTA for property-guided molecule generation in 3D while achieving much faster inference times. We perform a systematic study of various property embedding methods and identify the optimal approach by property type and property evaluation tasks. We achieve SOTA performance of property-guided generation across many properties while maintaining SOTA performance for standard metrics, such as molecule stability and PoseBusters validity. We propose a new task for property-guided generation that evaluates PropMolFlow models’ capability in generating out-of-distribution molecules. In a first-of-its-kind analysis, we perform extensive density functional theory-based (DFT) calculations to validate the accuracy of our property-guided generation. We highlight the importance of DFT calculations for different tasks and properties, and release a curated DFT dataset of 10,773 molecules generated by PropMolFlow for broader community use. We propose “closed-shell ratio” and revised atomic and molecular stability as key metrics based on valence-electron configuration to filter out open-shell molecules and molecules with invalid valency–charge configurations, because the QM9 training data consists exclusively of closed-shell and charge-neutral molecules.
Results
@@
Figure 1: Overview of the PropMolFlow methodology. a, PropMolFlow models are jointly trained on a molecular graph and property embedding. A molecular graph includes node scalar features, node position features and edge bond features. A property embedding comprises an optional Gaussian expansion mapping (G.E.) followed by a multilayer perceptron (MLP) that projects a scalar property to a high-dimensional embedding space. Conditioning on the property is achieved by the interaction between the property embedding and node scalar features in the molecular graph. b, A joint flow matching process is used for molecule generation by learning different molecule modalities together. c, Five possible interaction types between a property embedding and a molecular graph. Node scalar features $[A,C]$ interact with a property embedding $P=\phi_{\mathrm{prop}}(k)$ via various operations. An MLP transformation $\varphi_{\theta}$ is applied to convert the dimension back to that of the original $[A,C]$, where necessary. $\odot$’ represents an element-wise Hadamard product, and $\oplus$’ stands for a ‘Concatenate’ operation. d, Gaussian expansion as augmented property embedding. Curves in the top panel correspond to five Gaussian basis functions that are evenly spaced between the minimum and maximum gap values. Centers of Gaussian basis functions are marked by gray dashed lines, and the red dashed lines represent the three example input values. The bottom panels show the function values for each Gaussian for two example inputs. In a molecule ball-and-stick representation, white, gray, red, and blue indicate H, C, O, and N, respectively.
@@
PropMolFlow builds on top of the state-of-the-art generative process using flow matching. A flow model learns the molecule graph distribution through training a velocity field conditioned on observed molecule samples. As shown in Figure 1a, PropMolFlow uses a fully-connected graph to represent a molecule and an SE(3)-equivariant graph neural network (GNN) to parameterize the velocity field. Property-guided generation is achieved through the interaction of a property embedding with the node scalar features of a sampled molecular graph. The property embedding is constructed by first transforming the raw property value using a fixed (non-trainable) Gaussian basis expansion, followed by a shallow multilayer perceptron (MLP) that produces the final trainable embedding. This Gaussian expansion is optional, and its utility depends on the specific property type and evaluation metric. Once trained, the SE(3)-equivariant graph neural network defines a joint flow matching process that denoises the molecular graph by simultaneously updating all modalities—such as atom types, bond orders, and 3D coordinates—to generate valid molecules (Figure 1b). Details of the joint flow matching are provided in Methods.
We explored five specific operations, including ‘Concatenation’, ‘Sum’, ‘Multiply’, ‘Concatenate + Sum’, and ‘Concatenate + Multiply’, for combining property embeddings with node scalar features (Figure 1c). The dimension of node scalar features is preserved after interacting with the property embedding, providing a seamless integration into the original architecture. The non-trainable Gaussian expansion layer is used to enrich the property representations prior to being mapped to a property embedding via an MLP. Figure 1d illustrates an example of Gaussian expansion for the HOMO-LUMO gap. The top panel shows the five Gaussian basis functions and the bottom panel lists the function values for two example target properties, indicating the activation of each Gaussian basis function. We investigate the effectiveness of different property embedding strategies for six molecular properties in the QM9 dataset. The property guidance design space includes ten possible methods depending on whether the Gaussian expansion is used and which type of property embedding operation is employed. Details of the Gaussian expansion are provided in Methods.
We trained PropMolFlow on the QM9 data with explicit hydrogen atoms [28, 30]. The QM9 data contains 133,885 molecules, and each molecule contains up to 9 heavy atoms and consists of 3–29 atoms (average: 18 atoms), composed of up to 5 elements: C, N, O, F, and H. Molecules are fully optimized and molecular properties are obtained with DFT at the level of B3LYP/6-31G(2df,p). For property-guided generation, we considered six molecule properties: polarizability ($\alpha$), HOMO-LUMO gap, HOMO energy, LUMO energy, dipole moment ($\mu$), and heat capacity ($C_{v}$). To generate molecular graphs with explicit bond and charge information, we employed the QM9 SDF file distributed by DeepChem [30]. Upon inspection, we discovered numerous inconsistencies: invalid bond orders and non-zero net charges in roughly 30,000 molecules, despite QM9’s requirement for charge-neutral, closed-shell valency [28]. We corrected these discrepancies by reassigning bond orders or charges to enforce valency-charge consistency for the vast majority of entries and used the corrected data to train our PropMolFlow models. After the correction, we improved the structural validity of generated molecules significantly. Previous works that used bond orders overlooked this issue [23, 26]; instead, they modified evaluation metrics to tolerate chemically inconsistent bond–charge combinations, hence inflating molecule stability estimates contrary to QM9’s intended constraints. A detailed comparison between models using previous problematic data and the corrected data, together with the comparison between previous stability metrics and the revised ones, can be found in Supplemental Table LABEL:tab:data_and_metric_compare. A detailed description of our procedure and the alignment with the original QM9 XYZ data is provided in Supplemental Section LABEL:sec:qm9_data_fix. The revised QM9 SDF data is provided on ColabFit Exchange [31, 32].
PropMolFlow achieves state-of-the-art property-guided generation
We evaluated PropMolFlow based on its ability to accurately generate molecules with the target property, the validity of the generated molecules (i.e., whether they adhere to chemical rules), and the speed of generation. These assessments were first conducted under the in-distribution (ID) property-guided molecule generation task. The ID task involves sampling a range of property values from the same distribution as the training data. As such, it assesses the model’s capability to generate structures spanning a broad spectrum of target properties and has been widely adopted in prior studies of property-guided generative models [8, 21, 25, 27].
@@
In this task, a conditional generative model was first used to generate molecules conditioned on property values sampled from the same distribution as the training data (e.g., QM9). These sampled property values served as the input conditions. A separate property predictor/regressor was then applied to estimate the properties of the generated molecules. The predicted property values were compared to the original input conditions, and the mean absolute error (MAE) was used as the evaluation metric. A lower MAE indicates more accurate property conditioning and, therefore, better model performance. In PropMolFlow, we used an SE(3) GVP network for molecule generation and, correspondingly, trained a separate GVP regressor for each of the six molecular properties. For evaluation, we generated 10,000 molecules with input properties sampled from the property distribution of QM9 training data. Further details of the GVP regressors are provided in Methods.
We compared PropMolFlow with five SOTA property-guided generative models, all trained on the QM9 dataset. Among these models, EEGSDE (Equivariant Energy Guided Stochastic Differential Equations) employs an energy function to guide the diffusion for property-guided generation [27]; EquiFM is the first equivariant flow matching model for 3D molecule generation [21]; GeoLDM introduces the first latent diffusion model for molecular geometry generation [8]; Geometry-Complete Diffusion Model (GCDM) constructs an SE(3) equivariant geometry-complete diffusion model using geometry-complete perceptrons for molecule generation [25, 12]; and JODO utilizes a diffusion graph transformer for the joint generation of 2D molecular graphs and 3D geometries. Overall, PropMolFlow achieves competitive performance against the SOTA models on property-guided generation (Table 1). It should be noticed that EEGSDE, EquiFM, GeoLDM, and GCDM do not include bond orders in their molecular graphs while JODO incorporates both bond existence and bond orders. All five baseline conditional models omit atomic charges, in contrast to PropMolFlow, which can generate molecules with non-zero atomic charges (see Methods for more details of the baseline models). The “Random (Upper-bound)” intuitive baseline fully shuffles property values, thereby removing any correlation between molecular structures and properties, and serves as an upper bound on error. The “# Atoms” baseline uses only the number of atoms as a predictor for molecular properties, capturing coarse size-dependent trends. The “QM9 (Lower-bound)” baseline corresponds to a pretrained EGNN regressor provided by Hoogeboom et al. [7], trained directly on QM9 and serving as a lower bound on achievable error. Improvement over the “Random” baseline indicates that the conditional generation effectively integrates property information. Surpassing the “# Atoms” baseline suggests that the generative model captures structural features beyond simple atom count when generating new molecules.
@@
Table 1: Performance of PropMolFlow with respect to property alignment. Results for the baseline models use an equivariant graph neural network (EGNN) as the regressor, whereas PropMolFlow results use our pretrained SE(3) equivariant GVP regressor. Best results are bolded, whereas the second-best results are underlined. JODO results were reported on our sampled molecules using publicly available model checkpoints.
@@
PropMolFlow 1.31 391 254 315 0.620 0.626
Notably, PropMolFlow shows the lowest MAE for two properties—$\alpha$ and $\mu$—and comparable performance for $C_{v}$ and $\epsilon_{\rm{HOMO}}$ versus the SOTA JODO, and the second best performance for the other two properties $\epsilon_{\rm{LUMO}}$ and $\Delta\epsilon$. For instance, PropMolFlow achieves substantially lower MAEs compared to GCDM for $\epsilon_{\rm{LUMO}}$ (315 meV vs. 479 meV), and $\mu$ (0.620 Debye vs. 0.844 Debye), demonstrating a clear advantage in controllable molecular generation over prior diffusion- and flow-based approaches. The superior performance of PropMolFlow can be ascribed to the inclusion of bond order information and discrete treatment for atom types, charges and bond orders, which encodes structural information more accurately and hence offers a more robust structure–property mapping for generated molecules. It also benefits from the thorough exploration of property-dependent optimal embedding methods (Supplementary Table LABEL:tab:mae_range and Supplemental Section LABEL:sec:model_stats). For instance, the MAE variation for $\Delta\epsilon$ can vary by more than 40%, from 391 meV using a ‘Concatenate_Sum’ embedding to 551 meV using a ‘Concatenate’ embedding, both without Gaussian expansions.

### v4

@@ Abstract
Molecule generation is advancing rapidly in chemical discovery and drug design. Flow matching methods have recently set the state of the art (SOTA) in unconditional molecule generation, surpassing score-based diffusion models. However, diffusion models still lead in property-guided generation. In this work, we introduce PropMolFlow, an approach for property-guided molecule generation based on geometry-complete SE(3)-equivariant flow matching. Integrating five different property embedding methods with a Gaussian expansion of scalar properties, PropMolFlow achieves competitive performance against previous SOTA diffusion models in conditional molecule generation while maintaining high structural stability and validity. Additionally, it enables faster sampling speed with fewer time steps compared to baseline models. We highlight the importance of validating the properties of generated molecules through DFT calculations. Furthermore, we introduce a task to assess the model’s ability to propose molecules with underrepresented property values, assessing its capacity for out-of-distribution generalization.
@@ Introduction
Deep generative models are promising to accelerate chemical discovery by statistically sampling molecular structures, reducing reliance on costly physics-based simulations [sanchez-lengeling2018]. State-of-the-art (SOTA) 3D molecule generation primarily uses diffusion with equivariant graph neural networks (EGNN), enabling accurate sampling of molecular geometries. [hoogeboom2022, satorras2022, xu2023]. Flow matching is an emerging alternative to diffusion models [liu2022, lipman2023, albergo2023], offering flexibility over priors and probability paths while improving sampling efficiency across materials [hoellmer2025], proteins [campbell2024], and small molecules [dunn2024-2].
Seminal diffusion [hoogeboom2022, xu2023] and flow-matching models [song2023] treat molecules as point clouds with continuous representations for discrete molecular modalities, and parameterize the generative process via E(3)-EGNN. However, continuous encodings disrespect inherently discrete features, such as atom types, and E(3) models cannot capture chirality [dumitrescu2024]. Recent studies also found bond orders in molecular graphs to be important in improving validity of generated molecules [vignac2023, dunn2024-2]. FlowMol [dunn2024-2] achieves SOTA unconditional molecule generation, by using geometry-complete SE(3) generative processes, explicit bond-order modeling, and discrete flow matching for atom types, formal charges and bond orders, though its extension to property-guided generation remains unaddressed.
@@
Property-guided molecule generation aims to design molecules that satisfy target properties, often by concatenating property values with node features in molecular graphs [hoogeboom2022, xu2023, morehead2024] — a strategy mainly explored on the single-molecule QM9 data [ramakrishnan2014quantum, wu2018moleculenet]. Although effective, this approach may oversimplify how properties interact with molecular structures; for instance, Gebauer et al. [gebauer2022] suggest Gaussian expansions can extract property information more robustly. Systematic studies of property embedding methods are lacking, and optimal strategies may vary by properties. To validate properties of generated molecules, prior methods use separate property predictors trained on the fully relaxed QM9 molecules, which may struggle to make accurate predictions on non-relaxed, generated molecules. Moreover, property-guided generation is typically evaluated in an ‘in-distribution’ setting, where property values and atom counts sampled within the QM9 distribution are used to generate molecules [hoogeboom2022]. Chemical discovery requires generating molecules with underrepresented or out-of-distribution (OOD) property values.
We introduce PropMolFlow, a property-guided molecule generation framework that integrates various property embedding methods with an SE(3)-equivariant flow-matching process [dunn2024-2]. We evaluate PropMolFlow on QM9, demonstrating its competitive performance versus previous methods, while achieving faster inference. Besides, we propose an out-of-distribution generation task, and validate properties of generated molecules using extensive density functional theory (DFT) calculations. Finally, we introduce the close-shell ratio and revised stability metrics to complement existing metrics and ensure more comprehensive evaluations for molecule generation.
Results and Discussion
@@
Figure 1: Overview of the PropMolFlow methodology. a, PropMolFlow models are jointly trained on a molecular graph and property embedding. A molecular graph includes node scalar features, node position features and edge bond features. A property embedding comprises an optional Gaussian expansion mapping (G.E.) followed by a multilayer perceptron (MLP) that projects a scalar property to a high-dimensional embedding space. Conditioning on the property is achieved by the interaction between the property embedding and node scalar features. b, A joint flow matching process is used to generate different molecule modalities together. c, Five interaction types between a property embedding ‘P’ and a molecular graph. Node scalar features $A_{t},C_{t}]$ include atom type $A$ and formal charge $C$ at time $t$. An MLP transformation $\varphi_{\theta}$ is applied to convert the dimension back to that of the original $A_{t},C_{t}]$, where necessary. $\odot$’ represents an element-wise Hadamard product, and $\oplus$’ indicates a ‘Concatenate’ operation. $E_{t}$ and $X_{t}$ represent respective bond edge features and node position features. d, Gaussian expansion as augmented property embedding. Curves in the top panel correspond to five Gaussian basis functions that are evenly spaced between the minimum and maximum property values. Centers of Gaussians are marked by gray dashed lines, and the red solid lines represent two example inputs. The bottom panels show the function values for each Gaussian for two inputs. In a molecule configuration, white, gray, red, and blue indicate H, C, O, and N, respectively.
@@
PropMolFlow builds on top of the FlowMol architecture. [dunn2024-2] The model generates samples by integrating over NN-parameterized conditional velocity fields [lipman2023, albergo2023, liu2022]. An SE(3)-EGNN based on geometric vector perceptrons (GVP) [jing2021] is used and molecules are represented by fully-connected graphs (Figure 1a). The generation is achieved by interacting property embeddings and node scalar features. Property embeddings are constructed by a non-trainable Gaussian expansion, followed with a shallow multilayer perceptron (MLP). This Gaussian expansion is optional, and its utility depends on property types and evaluation tasks. Once trained, the NN defines a joint flow matching that denoises the molecular graph by simultaneously updating all modalities—such as atom types, bond orders, and 3D coordinates—to generate molecules (Figure 1b). Details of flow matching and neural networks are provided in the ‘Joint Flow Matching’ and ‘Model Architecture’ sections.
Five embedding methods and usage of Gaussian expansion are explored (Figure 1c,d). We evaluate the property-guided generation over six molecular properties in QM9. Details for property embeddings are provided in the ‘Property Embedding Operations’ and ‘Gaussian Expansion’ sections. PropMolFlow is trained on a revised QM9 SDF data with explicit hydrogen atoms, bond orders and formal charges [ramakrishnan2014quantum, wu2018moleculenet]. This revised SDF file fixed the bond and charge inconsistencies of the original SDF file distributed by DeepChem [wu2018moleculenet]. Data details are provided in the ‘Data’ section.
Competitive property-guided generation with PropMolFlow
@@
We evaluated PropMolFlow on its ability to generate molecules with target properties, molecular structural validity, and inference speeds. These assessments were conducted under the in-distribution (ID) task. This task assesses the model’s capability to generate structures spanning a spectrum of joint target properties and atom counts within data distributions and has been adopted in prior studies [xu2023, song2023, morehead2024, bao2023]. Generative models sample molecules by property inputs, and a separate property predictor evaluates properties of generated molecules. Predicted values are compared to input values, and the mean absolute error (MAE) was used as the evaluation metric. A lower MAE indicates better model performance. Further details of how to sample molecules and GVP regressors are provided in the ‘Inference’ and ‘GVP regressor details’ sections.
We compared PropMolFlow with five baseline models, including EEGSDE [bao2023], EquiFM [song2023], GeoLDM [xu2023], GCDM [morehead2024] and JODO [huang2024], all trained on QM9. Two intuitive baselines are also included to represent the respective upper and lower bounds of MAEs, namely “Random (Upper-bound)” and “QM9 (Lower-bound)”, along with another baseline “# Atoms” (see the ‘Baselines’ section for more details).
@@
Table 1: Performance of PropMolFlow with respect to property alignment. Results for the baseline models use an equivariant graph neural network (EGNN) as the regressor, whereas PropMolFlow results use our pretrained SE(3) GVP regressor as well as an EGNN regressor. JODO results were reported on our sampled molecules using publicly available model checkpoints.
@@
PropMolFlow, GVP 1.31 391 254 315 0.620 0.626
PropMolFlow, EGNN 1.36 391 246 312 0.632 0.640
Overall, PropMolFlow achieves competitive performance against the SOTA models (Table 1). It shows the lowest MAE for $\alpha$ and $\mu$, and comparable performance for $C_{v}$ and $\epsilon_{\rm{HOMO}}$ versus the SOTA JODO, and the second best performance for $\epsilon_{\rm{LUMO}}$ and $\Delta\epsilon$. The superior performance of PropMolFlow benefits from the thorough exploration of property-dependent optimal embedding methods (Supplementary Table LABEL:tab:mae_range and Supplementary Section LABEL:sec:model_stats). For instance, the MAE for $\Delta\epsilon$ can vary by more than 40%, from 391 meV using a ‘Concatenate_Sum’ embedding to 551 meV using a ‘Concatenate’ embedding, both without Gaussian expansions. To avoid bias of comparison using different predictors, we also report MAEs using EGNN predictors released by Bao et al.  [bao2023] in Table 1. It confirms that different predictors produce consistent MAEs.